l th steps from Q076 Uniform Probability Distribution see notes amount of time, in minutes, that a person must wait for a bus is uniformly distributed between zero and The 29 minutes, inclusive a. Draw the graph. (4 points) P(. 0 3448 29 24 b. Find the waiting time that corresponds to the 8oh percentile value. This is the waiting that a person will have to wait 80th of the time. 5 points K-80% * 2.9 K-0.80メ24 will Wa + 23.2 minutes go% of the time a Person
Answer #1

Solution :

Given that,

a = 0

b = 29

The 80th percentile is

= a + (b -a) p

= 0 + (29-0) * 0.80

= 29*0.80

= 23.2